OpenStudy (anonymous):

Kite DCFE is inscribed in circle A shown below. If the measure of is DEF 232°, what is the measure of ∡DEF?

OpenStudy (anonymous):

Wrong picture? No E and F points on it.

OpenStudy (anonymous):

my bad lol

OpenStudy (anonymous):

there ^^ hehehe i got it

OpenStudy (anonymous):

Arc DEF = 232 degrees. So, arc DCF = 360-232 = 128. Now, you use the theorem Directrix was using (wrongly) in the other problem. And, you get 128/2 = 64 degrees.

OpenStudy (anonymous):